A business man borrowed P10000 with interest at the rate of 5% payable annually. The debt will be paid, principal and interest included, by equal payments at the end of each year for 5 years. Compute the annual payment and construct the amortization schedule.
Answer #1

ANNUAL PAYMENT: 2,309.75

AMORTIZATION SCHEDULE:

Year Payment Principal Interest Balance
0                 -                   -               -      10,000.00
1    2,309.75    1,809.75    500.00      8,190.25
2    2,309.75    1,900.24    409.51      6,290.01
3    2,309.75    1,995.25    314.50      4,294.76
4    2,309.75    2,095.01    214.74      2,199.75
5    2,309.75    2,199.76    109.99            (0.01)

*(0.01) is rounding difference.

  • The loan is repaid in equal installments. Therefore, these installments are in the nature of an annuity.
  • The present value of loan = 10,000 (i.e. present value of annuity)
  • Present value of annuity = Annuity × PVAF (i%, n years)
    • i.e. Annuity = Present value of annuity ÷ PVAF (i%, n years)
    • Annuity = 10,000 ÷ 4.32948 = 2,309.75
  • This annual payment includes principal repayment and interest calculated on ending balance (of the previous year which is outstanding on the last day of the year; principal is repaid at the end of each year for 5 years).
    • E.g. Interest included in the first installment = 10,000 × 5% = 500
    • The remaining portion of the installment is principal repayment. Therefore, principal repayment included in the installment = annual installment - interest as calculated above. E.g. Principal repayment included in the first installment = 2,309.75 - 500 = 1,809.75
  • Balance: It is the balance outstanding at the end of the year. Balance at the end of the year = Previous ending balance - principal repayment during the year.
  • PVAF: Present value of annuity factor can be calculated using the following formula:
  • PV PVAF = 1 -(1+r) Y =
    • r = interest rate
    • n = number of periods
      • PVAF(5%, 5 years) = (1 - (1+0.05)-5 )÷ 0.05 = 4.32948